OpenStudy (anonymous):

solve by using the square root property. (2k-5)^2=16

OpenStudy (tkhunny):

2k-5 = 4 or 2k-5 = -4 ?? What else would a "Square Root Property" be?

OpenStudy (anonymous):

So you don't do anything with the 2 and the -5

OpenStudy (anonymous):

not until you solve \[2k-5=4\] and \[2k-5=-4\] for \(k\)